### v1

# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

Abstract.
Answering natural language (NL) questions about tables, which is referred to as Tabular Question Answering (TQA), is important because it enables users to extract meaningful insights quickly and efficiently from structured data, bridging the gap between human language and machine-readable formats. Many of these tables originate from web sources or real-world scenarios, necessitating careful data preparation (or data prep for short) to ensure accurate answers. However, unlike traditional data prep, question-aware data prep introduces new requirements, which include tasks such as column augmentation and filtering for given questions, and question-aware value normalization or conversion. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Iteratively executes and debugs the generated code to ensure correct outcomes. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the SOTA TQA solutions through question-aware data preparation.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://anonymous.4open.science/r/AutoPrep-Ano.
1. Introduction
Tabular Question Answering (TQA) refers to the tasks of answering natural language (NL) questions based on provided data tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and thus has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
(a) WikiTQ
(b) TabFact
Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets, with statistics calculated from a sample of 500 instances from each dataset.
(a) Question-Aware Column Augmentation: Necessary information for QA is not explicitly available in tables.
(b) Question-Aware Value Normalization: Heterogeneous data representations or formats in tables, and how to normalize is query-dependent.
(c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
Figure 2. Examples of Question-Aware Data Preparation for TQA tasks.
Because TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024) naturally rely on large language models (LLMs). These solutions either utilize LLMs as black-boxes or leverage them for code generation. However, question-aware data preparation is not explicitly explored in these approaches, which limits the effectiveness of LLM-based methods.
Figure 1 presents the results of an error analysis of LLM-based code generation (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The analysis reveals that 84% and 76% of the errors stem from insufficient data prep at various granularities (i.e., table-level, column-level, and cell-level) in relation to the NL questions. A similar trend is observed when using LLMs as black-boxes, though those results are not displayed here due to space constraints. These findings underscore the necessity for meticulous data prep to ensure accurate answers.
To better understand these cases, let’s illustrate by examples.
Example 0.
Figure 2 shows several examples of question-aware data prep tasks that are needed for TQA.
(a) Question-Aware Column Augmentation. Figure 2(a) illustrates the necessity of augmenting the tables with new columns, as the information that NL questions require is not explicitly available. For example, to answer $Q_{1}$, a new Country column should be created to augment $T_{1}$, e.g., by extracting country information from Cyclist.
(b) Question-Aware Value Normalization: Figure 2(b) highlights the need for normalizing data types or formats in tables (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$). Note that normalizing tables in TQA requires converting data types or formats based on the need of questions, meaning different questions may result in various operations. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Question-Aware Column Filtering. Figure 2(c) highlights the importance of selecting question-relevant columns from tables, as irrelevant columns add complexity and can lead to errors in code generation. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may mislead the LLM. Additionally, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$. Providing all columns to the LLM can thus result in long-context understanding challenges (Li et al., 2024).
Example 1.1 highlights that different types of data prep operations are needed to satisfy the need of different NL questions.
Given that data prep is a complex challenge without a one-size-fits-all solution, a sensible approach is thus to develop specialized methods w.r.t. each type of data prep operations. These specialized methods, such as data normalization, can involve either LLM-powered code generation or retrieving existing code from data prep libraries. An intelligent module can then automatically determine which operations to perform based on the properties of the question and the corresponding table.
AutoPrep: A Multi-Agent LLM Framework. To address the problem, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables based on the specific needs of NL questions, eliminating the need for users to specify data prep operations. The core idea is to decompose the data prep task in TQA into sub-tasks, leveraging multiple agents specialized in different areas to ensure accurate and contextually relevant responses.
Given an NL question over a table, AutoPrep performs data prep through the following three core components:
  • Planner: This component is responsible for analyzing the natural language question and the structure of the given table. It generates a logical plan that outlines the sequence of high-level operations required to prepare the data for answering the question. These operations may include tasks such as column augmentation, value normalization, or filtering irrelevant data.
  • Programmer: Once the logical plan is in place, the Programmer translates it into a physical plan by generating low-level executable code. This component is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats.
  • Executor: The Executor is tasked with running the generated code and ensuring its correct execution. It performs iterative execution, where each operation is applied sequentially to the table, and any issues or errors are detected and debugged in real-time.
Technical Challenges. Two key technical challenges arise in our proposed AutoPrep framework.
(C1) Planner: High-Level Operation Suggestion. Even for the same table, different NL questions may require not only distinct data prep operations but also varying sequences of those operations. Thus, suggesting a tailored sequence of high-level operations that meet the specific needs of these questions is a significant challenge.
(C2) Programmer: Low-Level Code Generation. A given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. Generating optimized code for each operation that is customized for the tables presents a non-trivial task. For instance, consider the string-to-integer conversion in Example 2.4; failing to account for different value formats can lead to execution errors from a simple type conversion.
Key Ideas. To address the challenge (C1), we propose a Chain-of-Clauses (CoC) based reasoning method for the Planner agent. This method analyzes the NL question and outputs a sequence of high-level desired data prep operations, namely an Analysis Sketch. Based on the Analysis Sketch, the Planner assigns high-level operations to specialized agents for low-level implementation (i.e., code generation).
To tackle challenge (C2), we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, design an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism.
Contributions. Our contributions are summarized as follows.
(1) We propose AutoPrep, an LLM-based multi-agent framework for automatic data prep in TQA. We formally define the problem and formalizes typical high-level data prep operations (Section 2), and introduce an overview of the AutoPrep framework (Section 3).
(2) We develop effective techniques for our AutoPrep framework. We propose a novel Chain-of-Clauses reasoning mechanism (Section 4) to suggest high-level operations and a tool-augmented method to generate low-level code (Section 5).
(3) We conduct a thorough evaluation on data prep in TQA (Section 6). Extensive experiments show that AutoPrep achieves new SOTA accuracy on well-adopted TQA datasets. We make the code used in our experiments available on an anonymous GitHub repo [^1] https://anonymous.4open.science/r/AutoPrep-Ano.
2. Preliminaries
This section first introduces Tabular Question Answering (TQA) in Section 2.1 and then formalizes the problem of question-aware data prep for TQA in Section 2.2. Then, we analyze the limitation of existing LLM-based methods in Section 2.3.
2.1. Tabular Question Answering
Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) ${A_{1},A_{2},\ldots,A_{m}}$ and $n$ rows ${r_{1},r_{2},\ldots,r_{n}}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table. The problem of tabular question answering (TQA) is to generate an answer, denoted as $A$, in response to question $Q$ based on the information in table $T$. By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
Example 0.
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
Note that TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to the specific needs of question $Q$.
Consider Example 2.1: careful data prep, such as extracting countries into a new column and then normalizing them as ISO codes, is crucial for generating an accurate answer. Moreover, it is important to note that different questions may require different data prep tasks, even for the same table. For instance, country extraction may not be necessary if the question is asking which cyclist, instead of country, won the most medals.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
  • Augment: a data prep task that augments a new column for table $T$ based on existing columns, to better answer $Q$. This task typically involves column combinations via arithmetic operations, value extraction and inference, etc.
  • Normalize: a data prep task that normalize types or formats of the values in a column of $T$ based on the needs of $Q$. This task typically involves value representation or format normalization, type conversion, etc.
  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answering question $Q$. This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$. For example, to implement the Augment operation $o_{1}$ that augments a Country column in $T_{1}$, a possible implementation $f_{i}$ is to write a customized program by using a specific regular expression as shown in Figure 2(a).
Example 0.
Figure 2 shows examples of question-aware data prep operations for TQA, along with their implementations in Python.
(a) The Augment operation: Figure 2(a) shows three examples of Augment operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
(b) The Normalization operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing country formats in $T_{4}$, converting strings into integers in $T_{5}$, and normalizing value formats in Date in $T_{6}$. We can see that, to implement a specific operation (e.g., in $T_{4}$), we sometimes need to call external tools, like a function clean_country from an external library.
(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{5}$ and $T_{6}$. The Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, potentially misleading the LLM. Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
Note that the above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, the set of data prep operations is extensible, in order to support new or domain-specific data prep tasks.
Question-Aware Data Prep for TQA. Given an NL question $Q$ posed over table $T$, the problem of question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan. Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
Example 0.
To better answer the example question $Q_{1}$ over $T_{1}$ in Figure 2(a), we first generate an Augment operation to transform $T_{1}$ into $T^{\prime}_{1}$ with three columns: Cyclist, Medal, and Country. Next, we generate a Normalize operation, similar to the one in $T_{4}$, to standardize the values in Country into a unified format, producing table $T^{\prime\prime}_{1}$. Finally, we leverage LLMs, either as black-box or for code generation, to compute the answer $A_{1}$ to $Q_{1}$ from the prepared table $T^{\prime\prime}_{1}$.
2.3. Limitations of Existing LLM-based Solutions
Naturally, LLMs can be utilized to support question-aware data prep for TQA. A straightforward way is to prompt an LLM with a task description and several demonstrations (i.e., examples), and requires the LLM to generate data prep code, as shown in Figure 3.
However, as shown in Figure 2, the diversity of NL questions necessitates distinct data prep operations. Next, let’s see how a single LLM is used to solve this problem, by an example.
Example 0.
Consider the example table in Figure 3 with columns Date, Cyclist, Medal and Age, along with question: “Which country has the most medals in total in February”. When using few-shot prompting with a task description and a few demonstrations, this strategy may face limitations in both stages of data prep in TQA.
First, as the number of demonstrations is limited, LLMs may struggle to decide which data prep operations are needed for the NL questions, leading to both false negatives (FNs) and false positives (FPs). For example, numeralizing Age in code block ② is an FP, as it is irrelevant to the question. Conversely, the LLM fails to normalize the Date column before extracting the month in code block ③, which is an FN.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), a single LLM may struggle to generate fine-grained code for each operation. For instance, normalizing Medal in code block ③ does not handle all edge cases (e.g., ‘2*’), and the country extraction in code block ⑤ fails to correctly process ‘Dav.ITA’, which is formatted differently from other values.
Example 2.4 shows that, because each of the above tasks is unique, a single model may not perform well across all scenarios. Thus, given that data prep is a complex challenge without a one-size-fits-all solution, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly.
Figure 3. An LLM-based solution that prompts an LLM with a task description and several demonstrations (i.e., exemplars), and requires the LLM to generate data prep code.
3. An Overview of AutoPrep
Figure 4. Our proposed Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA method (Right), such as NL2SQL as shown in the Figure. The proposed framework AutoPrep decomposes the data prep process into two stages: ① The Planning stage and ② the Executing stage.
To address the limitations, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables for given NL questions. Figure 4 provides an overview of our framework. Given an NL question $Q$ posed over table $T$, AutoPrep decomposes the data prep process into three stages:
  1. (1)
    [Uncaptioned image] Planner Agent: the Planning stage. It guides the LLM to suggest high-level data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
  2. (2)
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
  3. (3)
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and debugs iteratively to improve the generated code quality.
After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table. This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper. For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4. Note that other strategies could also be used by the agent in a “plug-and-play” manner, which will be discussed in the experiments.
Example 0.
Figure 4 illustrates how AutoPrep supports data prep for question “Which country has the most medals in total in February?” posed over a table with columns Date, Cyclist, Medal and Age.
(a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
  • $T^{\prime}=$Filter($T$, [Date, Cyclist, Medal]) that filters out column Age, which is irrelevant to the question;
  • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires aggregating the “medals in total”;
  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively. Each specialized Programmer focuses on generating executable code for its assigned operations.
(c) The Executing stage: a Executor agent iteratively refines the generated code if any errors occur.
After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
The Planner agent and the three Programmer agents (i.e., Augmenter, Normalizer, and Filter) are the core modules of AutoPrep, which will be discussed in more details below.
The Planner Agent. Given a question $Q$ and a table, the Planner agent generates a sequence of high-level operations for the specific question, as illustrated in the previous example. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. Given a high-level operation $o_{i}$, a Programmer agent is responsible for generating low-level code $f_{i}$ to execute the operation. As discussed previous, we develop specialized agents i.e., Augmenter, Normalizer, and Filter for various types of operations. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
Note that our proposed AutoPrep framework is extensible. When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them. The central Planner agent can then determine which operations should be performed and assign them accordingly.
4. The Planner Agent
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question. In this section, we first introduce a direct prompting method in Section 4.1 that better demonstrates the key challenge in designing the Planner agent. Then, to address the challenge, we propose our Chain-of-Clauses (CoC) reasoning method in Section 4.2.
Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach, as shown in Algorithm 1. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
Algorithm 1 Direct Prompting for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:// Prepare $k$ in-context exemplars (i.e., demos)
3:$\mathcal{D}=\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$
4:// Prompt the LLM for inferring the operations
5:$O\leftarrow\theta(T,Q,\Sigma,\mathcal{D})$
6:return $O$
Specifically, Algorithm 1 first prepares a set $\mathcal{D}$ of $k$ exemplars $\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$ that demonstrates to LLM $\theta$ the goal of this logical plan generation task. Then, the algorithm prompts the LLM to infer the operations $O$ for our input table $T$ and question $Q$.
Example 0.
Figure 5(a) shows an example of the aforementioned direct prompting method, which produces two high-level operations, Filter and Normalize. However, this logical plan might not be accurate as discussed as follows.
(a) Incorrect operations: The Filter operation incorrectly filters out the Cyclist column because the column is not explicitly mentioned in the question. This operation is clearly disastrous, as the country information is implicitly embedded in the Cyclist column.
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Augment operation on Cyclist is not generated, as the column has already been filtered out. Additionally, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
Key Challenge. The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$. For instance, as shown in Figure 4, the Normalize operation on the Date column is generated by considering two key factors: (1) the question contains “in February”, and (2) the values in the Date column are inconsistent.
4.2. A Chain-of-Clauses Method
To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method the decomposes the entire process of logical plan generation into two phrases, as shown in Figure 5(b).
  • The first phase leverages an LLM to generate an SQL-like Analysis Sketch, outlining how the table should be transformed to produce the answer.
  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., Date LIKE02-%’. For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation due to the following reasons. First, our method decomposes the question into a set of analysis steps over the table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation. More importantly, our method jointly considers each analysis step along with the corresponding relevant data to prompt the LLM, effectively capturing the relationships between the question and the table.
Algorithm 2 CoC Reasoning for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:$O\leftarrow\emptyset$
3:// Phase I: Analysis Sketch generation
4:$S\leftarrow\theta(Q,T,\{(Q_{i},T_{i},s_{i})\})$
5:// Phase II: High-level operation generation
6:for each clause $s$ in $S$ do
7:    $T^{s}\leftarrow$ extract_data($T$)
8:    $o\leftarrow\theta(s,T^{s},\Sigma,\{s_{i},T^{s}_{i},\Sigma,o_{i}\})$
9:    $O\leftarrow O\cup\{o\}$
10:end for
11:return $O$
Specifically, Algorithm 2 presents the pseudo-code of our CoC reasoning method, which takes as input a question $Q$, a table $T$, a specification set $\Sigma$, and an LLM $\theta$ to produce an operation set $O$. To this end, we introduce a key concept of SQL-like Analysis Sketch, which is formally defined as follows.
SQL-like Analysis Sketch. An Analysis Sketch $S$ is an SQL-like statement, which can be represented as follows.
SELECT $A~{}|~{}{\tt agg}(A)~{}|~{}f(\{A_{i}\})$ FROM $T$
WHERE ${\tt Pred}(A_{i})$ AND $\ldots$ AND ${\tt Pred}(A_{j})$
GROUP BY $A$ ORDER BY $A$ LIMIT $n$
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f(T)$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
Given the key concept Analysis Sketch, Algorithm 2 produces a logical plan via two phases, as shown in the following example.
Example 0.
Figure 5(b) shows our proposed CoC reasoning method for the example question and table in Figure 4. Specifically, the method generate a set $O$ of operations via the following two phases.
(a) Phase I - Analysis Sketch Generation: In this phase, the algorithm prompts the LLM $\theta$ with several exemplars $\{(Q_{i},T_{i},s_{i})\}$ to generate an Analysis Sketch $S$, as shown in Figure 5(b).
(b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
  • $f(T)$ AS Country: this clause and the entire table $T$ are used to prompt the LLM to generate an Augment operation.
  • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize operation.
Finally, the algorithm prompt the LLM to generate a Filter operation that only selects the columns relevant to the Analysis Sketch.
5. The Programmer & Executor Agents
Figure 6. Our proposed tool-augmented method for physical plan generation and execution.
The Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for iterative code execution and debugging. A straightforward approach is to prompt an LLM with the logical plan using typical in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan. The code is then executed iteratively, and if any runtime errors occur, the Debugger is prompted with the error messages for debugging.
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper. Specifically, we first introduce the key idea of our tool-augmented method in  5.1, then present the function pool in Section 5.2, and finally develop an algorithm for code generation and execution in Section 5.3.
5.1. Tool-Augmented Method: Key Idea
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O={o_{1},\ldots,o_{|O|}}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
Function Selection and Argument Inference. The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Aug}$) and infer the arguments (e.g., regular expression) for the selected function.
For instance, given the Augment operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Aug}$ designed specifically for Augment, obtaining an extract function with two arguments: column and func. The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation if no existing function is suitable to meet the operation’s requirements.
Code Execution and Debugging. Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$ and support iterative debugging if there are any runtime errors or incorrect results.
For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values. In this case, based on the execution results, the Executor agent will refine the regular expression in the lambda function until no runtime errors or incorrect results occur.
After completing these two steps, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., the Normalize operation, as shown in Figure 6.
5.2. Search Space of Function Pools
This section presents a search of function pools for the different types of high-level operations formalized in the current AutoPrep.
Function pool for $\mathtt{Augment}$. We define a specific Programmer agent, i.e., Augmenter, to generate low-level code for the operation type $\mathtt{Augment}$ from the following function pool $\mathcal{F}_{\mathtt{Aug}}$.
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to the lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
  • $\mathtt{calculate}$. This function generates a new column through arithmetic operations of columns such as addition, subtraction, multiplication, or a combination of these. To use this operator, we need to provide two arguments: ${\tt columns}$ and ${\tt func}$. The argument ${\tt columns}$ should be a list of source column names. Similar to $\mathtt{extract}$, the ${\tt func}$ argument is typically a lambda function. However, for $\mathtt{calculate}$, the input to the function should be a dictionary of values rather than a string, since we may perform calculations on multiple columns. For example, given a table with two columns A and B, representing the scores of team A and team B, we need to generate a new column ScoreDifference. In this case, we set ${\tt columns}$ to [’A’, ’B’] and ${\tt func}$ to the specific lambda function lambda x: x[’A’]-x[’B’].
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates new strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
  • $\mathtt{infer}$. This function directly uses LLMs to deduce values that could not be processed by the functions above. It takes a list of source columns and the failed target values as input, learning from demonstrations of successfully processed values to directly output the target values.
Function pool for $\mathtt{Normalize}$. We define a specific Programmer agent, i.e., Normalizer, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ set to Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
  • $\mathtt{format\mbox{-}datetime}$. The function standardizes the format of columns with DateTime values. It has two arguments: ${\tt column}$ and ${\tt format}$. The ${\tt format}$ argument specifies the desired format for standardization. Given the table $T_{6}$ in Figure 2(b), to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%m-%d”.
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans_{d}ict}$. The $\tt trans_{d}ict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $trans_{d}ict$ to {’Italia’: ’ITA’}.
  • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
Function pool for $\mathtt{Filter}$. We design the Normalizer agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$. This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
Discussions. It is important to note that the search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools. Moreover, AutoPrep can also include external API functions, such as specific data prep functions for various semantic types (e.g., date, name, company) developed in the dataprep library (Peng et al., 2021).
5.3. Code Generation and Execution
Algorithm 3 Low-level Code Generation and Execution
1:A question $Q$, a table $T$, a number $r$, a high-level operation set $O$ and three function pools $\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$ and $\mathcal{F}_{\tt Filter}$, an LLM $\theta$
2:$F\leftarrow\varnothing,~{}T_{1}\leftarrow T$
3:for each operation $o_{i}$ in $O$ do
4:    // Step 1: Function selection and argument inference
5:    $\mathcal{F}\leftarrow$ Assign($o_{i}$, [$\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$, $\mathcal{F}_{\tt Filter}$])
6:    $f_{i}\leftarrow\theta(o_{i},Q,T,\mathcal{F})$
7:    // Step 2: Code generation and execution
8:    while $f_{i}$ has errors &\&& ${\tt debug\_round}<r$ do
9:         $f_{i}\leftarrow$ Debug($Q,T,f_{i}$)
[…]
11:end for
12:return $T_{|O|+1}$
Algorithm 3 presents the pseudo-code of our tool-augmented approach for code generation and execution. The algorithm takes as input a question $Q$, a table $T$, a set $O$ of high-level operations, three function pools, and an LLM, and produces a prepared table $T_{|O|+1}$. The algorithm iteratively examines the operations in $O$. In each iteration, it first generates a function from the corresponding function pool and refines it through multiple rounds of debugging. The algorithm then applies the generated function to the corresponding intermediate table and proceeds to the next iterations.
6. Experiments
6.1. Experimental Setup
Table 1. Statistics of Datasets.
Dataset # Num. Row. Col. Ans.
WikiTQ $22,033$ 4$\sim$753 3$\sim$25 string / list (3.05%)
TabFact $2,024$ 5$\sim$47 5$\sim$14 true / false (49.60%)
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019). The statistics are shown in Table 1.
  • WikiTQ. WikiTableQuestions (WikiTQ) is a dataset designed for QA tasks, containing complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
  • TabFact. Table-Fact-Checking (TabFact) dataset provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Khot et al., 2022) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the gold answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer pairs like 7.45 and 7.4499, 4 years and 4, etc. Afterwards, we adopt accuracy as our evaluation metric.
Baselines. We evaluate AutoPrep using two categories of methods: (a) TQA w/o Data Prep: Zero-Shot QA and Few-Shot QA (Chen, 2022), as well as Text-to-SQL (Rajkumar et al., 2022); (b) TQA with Data Prep: Binder (Cheng et al., 2023) and ReAcTable (Zhang et al., 2024). We further describe them below.
  • Zero-Shot and Few-Shot QA (Chen, 2022). Zero-Shot QA directly asks an LLM to generate the answer based on the serialized table and question. To improve the performance, Few-Shot QA further provides several demonstrations to enhance the In Context Learning ability of the LLM. We implement these two methods with the prompt provided by (Cheng et al., 2023).
  • Text-to-SQL (Rajkumar et al., 2022). Text-to-SQL is a natural code generation method for the TQA tasks. Consider a serialized table and a question, it first generates the SQL program and then executes it to get the final answer from the table for the question.
  • Binder (Cheng et al., 2023). Binder further enhances the Text-to-SQL method by integrating LLMs into SQL programs. It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone. We utilize the code provided by (cod, 2023).
  • ReAcTable (Zhang et al., 2024). ReAcTable uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs. Once all relevant data is gathered, it asks the LLMs to predict the answer. We run the original code from (cod, 2024) and kept all settings as default.
Implementation Details. We mainly use DeepSeek (Guo et al., 2024) API [^2] https://platform.deepseek.com/usage model as the foundation model for each LLM-based agent. We also explore other proprietary LLMs, such as GPT-3.5 (Brown et al., 2020), using the GPT-3.5-turbo-0613 version, as adopted by (Wang et al., 2024b). To ensure reproducibility, we set the temperature parameter for all agents to 0.01.
6.2. Data Prep Method Comparison
Table 2. Comparison of AutoPrep with data prep methods for TQA (the best results are bolded, and the second-best results are underlined).
Method DeepSeek GPT3.5
WikiTQ TabFact WikiTQ TabFact
Binder $56.98$ $79.64$ $\underline{56.74}$ $\underline{79.17}$
ReAcTable $\underline{64.25}$- $51.80$ $72.80$
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$ (87.01) $\mathbf{63.86}$ $\mathbf{84.29}$
Exp-1: Comparison of AutoPrep with other data prep methods for TQA. We conduct experiments to evaluate AutoPrep against current state-of-the-art data prep methods for TQA on two datasets: WikiTQ and TabFact. Specifically, we compare AutoPrep with Binder and ReAcTable, presenting the results in Table 2. Notably, due to the lack of public scripts for TabFact from ReAcTable (Zhang et al., 2024), we report results only for DeepSeek on WikiTQ and use GPT-3.5 results from their original paper.
The results clearly demonstrate that AutoPrep achieves state-of-the-art performance on both the WikiTQ and TabFact datasets when using DeepSeek and GPT-3.5 as the foundation model. In particular, AutoPrep outperforms Binder by 14.86% on WikiTQ and 8.50% on TabFact with DeepSeek. When using GPT-3.5, the performance improvements are 12.55% and 6.47%, respectively. Furthermore, when compared with ReAcTable, AutoPrep enhances performance by 1.87% on WikiTQ using DeepSeek. Remarkably, with GPT-3.5, AutoPrep surpasses ReAcTable by 23.28% and 15.78% on WikiTQ and TabFact, respectively.
These improvements are attributed to AutoPrep’s comprehensive data prep operations, which are not fully addressed by Binder and ReAcTable. While Binder integrates LLMs to generate new columns to supplement external knowledge, it neglects critical aspects such as column filering and value normalization. In contrast, ReAcTable employs the ReAct data processing paradigm but lacks a specific decomposition of data prep tasks. Our experimental results highlight the importance of thorough data prep in enhancing TQA performance, which is often a significant bottleneck for existing methods.
Additionally, as shown in Table 2, DeepSeek consistently outperforms GPT-3.5 while incurring lower API costs [^3] https://platform.deepseek.com/api-docs/zh-cn/quick_start/pricing/ https://openai.com/api/pricing/. Therefore, we have selected DeepSeek as our default foundation model for subsequent experiments.
6.3. Improvement of Data Prep for TQA
Table 3. Impact of data prep on TQA Methods.
Method  WikiTQ  TabFact
Zero-Shot QA $17.66$ $76.78$
    + AutoPrep $28.22_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 10.56}}$ $78.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 1.92}}$
Few-Shot QA $56.05$ $81.97$
    + AutoPrep $56.33_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.28}}$ $82.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.73}}$
Text-to-SQL $51.66$ $68.53$
    + AutoPrep $\mathbf{65.45_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 13.79}}}$ $\mathbf{86.41_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 17.88}}}$
(a) Zero-Shot QA.
(b) Few-Shot QA.
(c) Text-to-SQL.
Figure 7. Performance on Small, Medium, Large tables from WikiTQ without/with data prep of AutoPrep.
Exp-2: Impact of data prep on TQA performance. We evaluate the performance of three baseline TQA methods, namely Zero-Shot QA, Few-Shot QA, and Text-to-SQL with and without the data prep provided by AutoPrep. The results are detailed in Table 3.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, Text-to-SQL experiences the most substantial enhancement, with performance increases of 26.69% on WikiTQ and 26.09% on TabFact. This significant improvement is attributed to Text-to-SQL’s sensitivity to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, Zero-Shot and Few-Shot QA methods also show considerable performance gains. Zero-Shot QA sees an improvement of 59.80% on WikiTQ and 2.50% on TabFact, while Few-Shot QA improves by 0.50% on WikiTQ and 0.89% on TabFact. These improvements are largely due to AutoPrep’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
To further validate the impact of data prep, we analyze performance across tables of varying sizes. We categorize the tables in WikiTQ and TabFact into small (fewer than 2048 tokens), medium (2048 to 4096 tokens), and large (more than 4096 tokens) categories. While all tables in TabFact are small, WikiTQ includes 4040 small tables, 200 medium tables, and 104 large tables. As shown in Figure 7, AutoPrep achieves significant gains of 5.88% and 23.67% on medium and large tables, respectively. In addition, Few-Shot QA demonstrates limited improvement for small tables, because it has already achieved a high accuracy.
These experimental results clearly show that the integration of AutoPrep enhances the performance of all TQA methods, which demonstrates the critical role of question-aware data preparation in optimizing TQA performance.
6.4. Evaluation on Planner Agent
Table 4. Comparison of different Methods for Planner and Programmer.
Agent Method  WikiTQ  TabFact
Planner Direct Prompting $58.17$ $83.35$
Chain-of-Clauses $65.45$ $86.41$
Programmer Code-generation $62.82$ $81.97$
Tool-augmented $65.45$ $86.41$
Exp-3: How does our proposed Analysis Sketch method in suggesting high-level operations compared with Direct Prompting method? We change the high-level operation suggestion methods in Planner and evaluate the framework variants on WikiTQ and TabFact dataset. The results are reported in Table 4.
As shown in the table, Analysis Sketch based method outperforms Direct Prompting method by 12.52% and 3.67%, which indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is much more significant than that in TabFact. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using direct prompting. In Summary, the experimental results prove that our proposed Analysis Sketch method is more effective in suggesting high-level operations, especially for difficult TQA instances.
6.5. Evaluation on Programmer Agents
Exp-4: How does our proposed tool-augmented method in generating low-level operations compared with existing code generation methods? We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default. We evaluate the performance on WikiTQ and TabFact and report the results in Table 4.
As illustrated in Table 4, our proposed tool-augmented method achieves better performance compared with the existing code generation method by an improvement of 4.19% and 5.42%. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. The above experimental results show the superiority of our method in both accuracy and cost.
Table 5. The average debugging round for Code-generation (Code) and Tool-augmented (Tool) Method.
Method  WikiTQ  TabFact
Code $1.35$ $1.28$
Tool $1.0$ $1.0$
Moreover, to compare the debugging method of these two methods, we record the average debugging rounds in Table 5. Our tool-augmented method conduct debugging process with fewer rounds, implying the effectiveness of our proposed debugging method for tool-augmented method.
6.6. Ablation Study
Table 6. Ablation Study of AutoPrep.
Method  WikiTQ  TabFact
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$
- Filter $62.78$ $84.98$
  - Augmenter $61.79$ $85.67$
  - Normalizer $62.02$ $83.79$
Exp-5: What is the contribution of each agent in AutoPrep? We conduct ablation experiments to evaluate the contribution of each agent in AutoPrep. Specifically, we ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
The results are reported in Table 6. For WikiTQ datasets, without column augmentation, the performance drops the most (5.92%). For TabFact datasets, the Normalizer matters the most with a performance drop by 3.13%. This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while column type normalization is a primary issue in TabFact datasets. Moreover, for both datasets, each agent plays an essential role in data preparation for TQA, which brings performance improvement by at least 4.25% and 0.86%.
The experimental results have demonstrated the rationality of the multi-agent design in AutoPrep. By solving three types of data preparation tasks in TQA methods in a specialized manner, the performance will be improved.
6.7. Case Study
(a) Case one
(b) Case two
Figure 8. Two case studies selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
This section presents two illustrative examples selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
As shown in Figure 8(a), to answer the question based on the original table, a Text-to-SQL method tries to filter all records with a score greater than 30. However, due to the limitation of SQL grammar, it can only use the operation Like to extract records where number 31, 32, …, or 39 is in values of column Score. This could miss out values with score greater than 40. Thus, it outputs error answer “3”. AutoPrep improves the performance of TQA methods by solving question-related data prep issues of the table in two stages. Specifically, based on the question in Figure 8(a), we first suggest two required high-level operations, i.e., a $\mathtt{Filter}$ to select the related column Score and a $\mathtt{Augment}$ to generate a new column from Score representing whether the team score is over 30. Then, we select appropriate functions from specific function pool and implement corresponding arguments to implement these high-level operations. Finally, we output a prepared table which only has columns related to the question. By passing the processed table and original question into the Text-to-SQL method, it generate a SQL that can easily extract the correct answer “4”.
Figure 8(b) illustrates a TQA instance with redundancy and inconsistency issues. To answer the question, it requires to first extract records with AirDate previous to “November 15, 2007” and then select records with RankOverall between 30 to 39. However, the values in column AirDate can not be directly assessed for chronological order through string comparison and RankOverall values are not in numerical type. Thus, although the generate SQL query is logically correct, it produces an wrong answer. AutoPrep suggests three high-level operations to solve this. Specifically, AutoPrep adopts an $\mathtt{Filter}$ to extract all question-related columns. And utilize two $\mathtt{Normalize}$ operations to format values of AirDate into sortable strings and cast RankOverall column to the integer type. We select three low-level operations to implement these high-level operations and get a new table with AirDate in %Y-%m-%d and RankOverall in integers. Then, we adopt the Text-to-SQL method on the prepared table, which generates a correct answer “3”.
7. Related Work
Tabular Question Answering. Most of the SOTA solutions for TQA rely on Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables. The existing solutions can be broadly divided into two categories: end-to-end reasoning (Wei et al., 2022) and code generation (Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024). End-to-end reasoning approaches, which prompt an LLM to directly generate answers through next token prediction (Wei et al., 2022), may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024). On the other hand, code generation approaches leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and achieving higher accuracy on real-world TQA tasks (Cheng et al., 2023; Zhang et al., 2024). For example, Dater (Ye et al., 2023) proposes to decompose questions and use code to deconstruct table contexts, enhancing end-to-end table reasoning with decomposed sub-tables and sub-questions. ReAcTable (Zhang et al., 2024) utilizes the ReAct paradigm (Yao et al., 2023) to generate codes which are executed step-by-step to extract intermediate question-related sub-tables for deducing the answer.
However, current LLM-based TQA methods mainly focus on the reasoning process of answer extraction instead of thorough and comprehensive data prep. While Binder (Cheng et al., 2023) proposes integrating SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks, it does not account for a wide range of data prep tasks, such as normalization and column filtering. In this paper, we highlight the significance of question-aware data prep for TQA which could be a bottleneck restricting the performance. We also point out primary challenges in question-aware TQA data prep and propose an automatic multi-agent framework to solve them.
Traditional Data Preparation. Data Prep techniques have been widely adapted to a wide range of tasks, such as machine learning (ML) and exploratory data analysis (EDA), as they turn raw data into a format that is ready for downstream tasks. A complete data preparation process usually involves a series of steps, such as data transformation, cleaning, and feature engineering. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify optimized data prep operation plan. Other approaches, such as Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020), apply meta-learning techniques to discover promising data prep configurations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) have explored the use of reinforcement learning to generate plans autonomously. For human-in-the-loop data prep, systems like DORIAN (Redyuk et al., 2022), Auto-Suggest (Yan and He, 2020), and Auto-Pipeline (Yang et al., 2021) incorporate recommendation-based methods to assist users. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
In this paper, we propose AutoPrep to generate question-specific data prep operations for TQA tasks. The key difference between these methods and AutoPrep is that the data prep operations generated by AutoPrep are tailored to specific user questions and customized for the tables, as different questions may have different data prep requirements, even for the same table.
LLMs-based Multi-Agent Framework. A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a). Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024). AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Autonomous Tabular Question Answering, which decomposes complex TQA problems step-by-step and design agents to identify and correct shortcomings in the problem-solving process.
While AutoTQA improves table analysisx, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
8. Conclusion
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and iteratively debugs. We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Additionally, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ and TabFact demonstrate that: (1) AutoPrep achieves state-of-the-art performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the overall performance of TQA methods.