### v2

# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

Abstract.
Answering natural language (NL) questions about tables, known as Tabular Question Answering (TQA), is crucial because it allows users to quickly and efficiently extract meaningful insights from structured data, effectively bridging the gap between human language and machine-readable formats. Many of these tables are derived from web sources or real-world scenarios, which require meticulous data preparation (or data prep) to ensure accurate responses. However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation. This question-aware data preparation involves specific tasks such as column augmentation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Executes the generated code to process the table. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the state-of-the-art TQA solutions through question-aware data preparation.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/fmh1art/AutoPrep.
1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024b) rely on large language models (LLMs), either utilizing LLMs as black-boxes or leveraging them for code generation.
(a) WikiTQ
(b) TabFact
Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets, with statistics calculated from a sample of 500 instances from each dataset.
(a) Question-Aware Column Augmentation: Necessary information for QA is not explicitly available in tables.
(b) Question-Aware Value Normalization: Heterogeneous data representations or formats in tables, and how to normalize is query-dependent.
(c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
Figure 2. Examples of Question-Aware Data Preparation.
Question-Aware Data Preparation. As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers. Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The results show that 84% and 76% of the errors arise from insufficiently addressing the data problems at various granularities (table-level, column-level, and cell-level) in relation to the NL questions. This highlights the critical role of thorough data prep in ensuring accurate responses for TQA.
Figure 2 shows several typical data problems critical to the quality of TQA, which in turn motivate the need of question-aware data preparation.
Example 0.
(a) Column-Level: Missing Column. The example in Figure 2(a) shows the Country column that the NL question $Q_{1}$ requires is not explicitly available.
In this case, an appropriate data prep could address this problem by extracting country information from Cyclist, i.e., augmenting the tables with a new Country column. Clearly, this data augmentation is driven by the specific need of the individual query.
(b) Cell-Level: Inconsistent Data Types or Formats. Figure 2(b) highlights the inconsistency among the data in different columns (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$).
This essentially demands data normalization. This normalization has to be question-aware, because different questions may require different types of normalization. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Table-Level: Irrelevant Columns. As shown in Figure 2(c), some columns in the tables are irrelevant to the questions. For example, $Q_{7}$ does not need the record column in table $T_{7}$. On the other hand, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
Filtering out the columns irrelevant to a question is important to TQA, as these irrelevant columns add complexity, potentially misleading LLMs to generate wrong answers. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may confuse the LLM when answering $Q_{7}$. Moreover, this column filtering, avoiding feeding all columns to the LLM, effectively reduces LLM inference cost.
The above examples demonstrate the urgent need of data prep to ensure the quality of TQA. However, blindly applying the existing data prep techniques to address all possible data problems tends to be neither effective nor efficient due to overlooking the specific needs of different NL questions.
In this work, we thus propose to study a new problem, namely question-aware data preparation. Unlike the conventional data prep which prepares tables once at an offline stage to serve many incoming analytics queries (questions), question-aware data prep addresses question specific data problems at online query stage. Once receiving a question, it tailors the tables to the specific need of this question, e.g., performing tasks such as question-specific column augmentation and filtering, as well as question-aware value normalization or conversion.
Key Challenges. Building a system to support question-aware data prep is challenging because of the diversity of both the data problems the tables encounter and the questions. As illustrated in the above examples, different questions may demand different types of data prep, e.g., augmentation, normalization, or filtering. Furthermore, even if facing the same type of data prep problem, different questions may require different ways to handle it. For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers. Existing LLM-based TQA solutions do not explicitly consider the question-aware data prep problem, thus suffering from the data issues mentioned above. For a detailed analysis, please see Section 2.
AutoPrep: A Hierarchical, Multi-Stage Approach. To address these challenges, we propose AutoPrep, which features two key ideas. First, inspired by the design principles of modern DBMS, in particular the distinction between the logical operations and their physical implementations, AutoPrep separates the high level, logical data prep operations and the concrete data prep methods. That is, it introduces a planning stage to generate a logical plan for each question. The logical plan is composed of a sequence of high-level data prep operations needed by the corresponding questions, including column augmentation, value normalization, or filtering irrelevant data, as shown in Figure 2. AutoPrep then introduces another stage to map the selected logical operations to the corresponding physical operations. In this way, AutoPrep decomposes a data prep task into smaller sub-tasks which are easier to solve than the original complex task. Moreover, it makes AutoPrep extensible, allowing us to easily develop new specialized implementations w.r.t. each type of data prep operations, or supporting new operation types.
Second, unlike conventional DBMS, in AutoPrep the physical operators w.r.t. each logical operator are not developed beforehand. Instead, given a data prep operation in the logical plan, AutoPrep generates a physical implementation specialized to the question on the fly. Taking the question specific need into consideration, this customized implementation thus tends to meet the unique need of different questions.
Driven by the above insight, we design the system into a multi-stage architecture.
The Planning Stage: Unlike DBMS where the logical operators are already available beforehand via SQL queries, AutoPrep has to decide on the appropriate logical data prep operations and in turn form the logical plan by analyzing the NL question and the given table. This is done during the planning stage.
The Programming Stage: This stage translates the logical plan into a physical plan by generating low-level executable code, which is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats, etc.
The Executing Stage: This stage executes the generated code for each operation and returning any errors encountered to the Programming stage for debugging.
We implements this AutoPrep architecture with the popular LLM-based Multi-Agent framework, which uses multiple small, independent agents to collaboratively solve complex problems.
More specifically, we design a Planner agent. Corresponding to the Planning Stage, it suggests a tailored sequence of high-level operations that meet the specific needs of these questions, leveraging the semantics understanding and reasoning ability of LLMs. The key technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer and guiding the agent’s reasoning based on this sketch. Compared with the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the relationships between questions and tables.
AutoPrep also contains a set of Programmer agent, each of which synthesizes a question specific implementation w.r.t. one logical data prep operation. However, existing LLM-based code synthesis often produces overly generic code that fails to effectively address the heterogeneity challenges of tables. For example, values may have diverse syntactic formats (e.g., “19-Oct” and “9/14” in $T_{6}$) or or semantic representations (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code tailored to such variations. To address this, we develop a tool-augmented approach that improves the LLM’s code generation capabilities by utilizing predefined API functions. This approach effectively resolves the issue of overly generic code by allowing the LLM to leverage specialized functions tailored to variations in table values. Moreover, corresponding to the Executing stage, we design an Executor agent that executes the generated code to process the table.
Contributions. Our contributions are summarized as follows.
(1) We introduce a novel problem of question-aware data preparation for TQA, which is formally defined in Section 2.
(2) We propose AutoPrep, an LLM-based multi-agent framework for question-aware data prep (Section 3). We develop effective techniques in AutoPrep for the Planner agent (Section 4) and the Programmer agents (Section 5).
(3) We conduct a thorough evaluation on data prep in TQA (Section 6). Extensive experiments show that AutoPrep achieves new SOTA accuracy, outperforming existing TQA methods without data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, and surpassing TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact.
2. Question-Aware Data Prep for TQA
2.1. Tabular Question Answering
Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) ${A_{1},A_{2},\ldots,A_{m}}$ and $n$ rows ${r_{1},r_{2},\ldots,r_{n}}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table. The problem of tabular question answering (TQA) is to generate an answer, denoted as $A$, in response to question $Q$ based on the information in table $T$. By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
Example 0.
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to specific needs of question $Q$.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
  • Augment: a data prep task that augments a new column for table $T$ based on existing columns, to better answer $Q$. This task typically involves column combinations via arithmetic operations, value extraction and inference, etc.
  • Normalize: a data prep task that normalize types or formats of the values in a column of $T$ based on the needs of $Q$. This task typically involves value representation or format normalization, type conversion, etc.
  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answering question $Q$. This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$.
Example 0.
Figure 2 shows examples of question-aware data prep operations for TQA, along with their implementations in Python.
(a) The Augment operation: Figure 2(a) shows three examples of Augment operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
(b) The Normalization operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing country formats in $T_{4}$, converting strings into integers in $T_{5}$, and normalizing value formats in Date in $T_{6}$. We can see that, to implement a specific operation (e.g., in $T_{4}$), we sometimes need to call external tools, like a function clean_country from an external library.
(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{5}$ and $T_{6}$. The Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, potentially misleading the LLM. Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
The above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, due to the separation of logical operations and physical implementations in AutoPrep, the set of data prep operations can be easily extended to support new data prep tasks.
Question-Aware Data Prep for TQA. Given an NL question $Q$ posed over table $T$, question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan. Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
Example 0.
To better answer the example question $Q_{1}$ over $T_{1}$ in Figure 2(a), we first generate an Augment operation to transform $T_{1}$ into $T^{\prime}_{1}$ with three columns: Cyclist, Medal, and Country. Next, we generate a Normalize operation, similar to the one in $T_{4}$, to standardize the values in Country into a unified format, producing table $T^{\prime\prime}_{1}$. Finally, we leverage LLMs, either as black-box or for code generation, to compute the answer $A_{1}$ to $Q_{1}$ from the prepared table $T^{\prime\prime}_{1}$.
2.3. Limitations of Existing LLM-Based Solutions
Figure 3. An LLM-based method with few-shot prompting for question-aware data prep.
A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions. Specifically, given the input token constraints of LLMs, it is often infeasible to supply an entire table and request a fully prepared table as output. Therefore, a more effective strategy employs few-shot prompting to generate data prep programs. For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”. The few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
However, this LLM-based solution may encounter the following limitations when performing question-aware data prep for TQA.
First, given the difficulties of understanding both NL questions and tables, it is challenging to determine which data prep operations are specifically required to meet the needs of each NL question, thus often resulting in false negatives (FNs) and false positives (FPs). Take the results in Figure 3 as an example: converting Age to a numerical format in code block ② is an FP, as it is irrelevant to the question. Conversely, the method fails to normalize the Date column before extracting the month in code block ③, which constitutes an FN, because it does not recognize the inconsistent Date formats.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues. For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
Recent methods, such as CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct. However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them at various granularities (table-level, column-level, and cell-level), within a single LLM agent. Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
Figure 4. A Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA methods (Right).
3. An Overview of AutoPrep
To address the limitations, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables for given NL questions. Figure 4 provides an overview of our framework. Given an NL question $Q$ posed over table $T$, AutoPrep decomposes the data prep process into three stages:
  1. (1)
    [Uncaptioned image] Planner Agent: the Planning stage. It guides the LLM to suggest high-level data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
  2. (2)
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
  3. (3)
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and reports error messages if any code bugs occur.
After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table. This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper. For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4. Note that other strategies could also be used by the agent in a “plug-and-play” manner, which will be discussed in the experiments.
Example 0.
Figure 4 illustrates how AutoPrep supports data prep for question “Which country has the most medals in total in February?” posed over a table with columns Date, Cyclist, Medal and Age.
(a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
  • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires aggregating the “medals in total”;
  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
  • $T^{\prime}=$Filter($T$, [Date, Country, Medal]) that filters out column Age, which is irrelevant to the question;
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively. Each specialized Programmer focuses on generating executable code for its assigned operations.
(c) The Executing stage: an Executor agent iteratively refines the generated code if any error occurs.
After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
The Planner Agent. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
Note that our proposed AutoPrep framework is extensible. When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them. The central Planner agent can then determine which operations should be performed and assign them accordingly.
4. The Planner
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question.
4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
Example 0.
Figure 5(a) shows an example of the aforementioned direct prompting method, which produces two high-level operations, Filter and Normalize. However, this logical plan might not be accurate as discussed as follows.
(a) Incorrect operations: The Filter operation incorrectly filters out the Cyclist column because the column is not explicitly mentioned in the question. This operation is clearly disastrous, as the country information is implicitly embedded in the Cyclist column.
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Augment operation on Cyclist is not generated, as the column has already been filtered out. Additionally, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
Key Challenge. The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$. For instance, as shown in Figure 4, the Normalize operation on the Date column is generated by considering two key factors: (1) the question contains “in February”, and (2) the values in the Date column are inconsistent.
4.2. A Chain-of-Clauses Method
To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method that decomposes the entire process of logical plan generation into two phases, as shown in Figure 5(b).
  • The first phase leverages an LLM to generate an SQL-like Analysis Sketch, outlining how the table should be transformed to produce the answer.
  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., Date LIKE02-%’. For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation. First, our method decomposes the question into a set of analysis steps over table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation. More importantly, our method jointly considers each analysis step along with the corresponding relevant columns (instead of the whole table) to prompt the LLM, effectively capturing the relationships between the question and the data.
SQL-like Analysis Sketch. An Analysis Sketch $S$ is an SQL-like statement, which can be represented as follows.
SELECT $A~{}|~{}{\tt agg}(A)~{}|~{}f(\{A_{i}\})$ FROM $T$
WHERE ${\tt Pred}(A_{i})$ AND $\ldots$ AND ${\tt Pred}(A_{j})$
GROUP BY $A$ ORDER BY $A$ LIMIT $n$
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f({\tt Cyclist})$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
Example 0.
Figure 5(b) shows our proposed CoC reasoning method for the example question and table in Figure 4. Specifically, the method generate a set $O$ of operations via the following two phases.
(a) Phase I - Analysis Sketch Generation: In this phase, the algorithm prompts the LLM $\theta$ with several exemplars $\{(Q_{i},T_{i},s_{i})\}$ to generate an Analysis Sketch $S$, as shown in Figure 5(b).
(b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
  • $f({\tt Cyclist})$ AS Country: this clause and relevant columns are used to prompt the LLM to generate an Augment operation.
  • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch.
5. The Programmer & Executor
Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for code execution and interactive debugging. A straightforward approach is to prompt an LLM with the logical plan using in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan. The code is then executed iteratively, and if any runtime errors occur, the Programmer is prompted with the error messages for debugging.
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in Figure 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper.
5.1. Tool-Augmented Method: Key Idea
Figure 6. Our proposed tool-augmented method for physical plan generation and execution.
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
Function Selection and Argument Inference. The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Aug}$) and infer the arguments (e.g., regular expression) for the selected function.
For instance, given the Augment operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Aug}$ designed specifically for Augment, obtaining an extract function with two arguments: column and func. The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation if no existing function is suitable to meet the operation’s requirements.
Code Execution and Debugging. Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$. If any bugs occur during execution, it captures and summarizes error messages and returns to corresponding Programmer agent. For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values. In this case, based on the execution results, the Executor agent records the error log and examines relevant data to summarize reasons, which will be passed to the corresponding Programmer to modify the code. After the execution and debugging process, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., a Normalize operation in Figure 6.
5.2. Search Space of Function Pools
This section presents a search space of function pools for different types of high-level operations formalized in the current AutoPrep.
Function pool for $\mathtt{Augment}$. We define a specific Programmer agent, i.e., Augmenter, to generate low-level code for the operation type $\mathtt{Augment}$ from the following function pool $\mathcal{F}_{\mathtt{Aug}}$.
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to a lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
  • $\mathtt{calculate}$. The function generates a new column through arithmetic operations of existing columns. It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns. For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [’2012’, ’2013’] and ${\tt func}$ to a lambda function lambda x: (x[’2013’]-x[’2012’])/x[’2012’].
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
  • $\mathtt{infer}$. This function directly uses LLMs to deduce values that could not be processed by the functions above. It takes a list of source columns and the failed target values as input, learning from demonstrations of successfully processed values to directly output the target values.
Function pool for $\mathtt{Normalize}$. We define a specific Programmer agent, i.e., Normalizer, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ as Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
  • $\mathtt{format\mbox{-}datetime}$. The function standardizes the format of columns with DateTime values. It has two arguments: ${\tt column}$ and ${\tt format}$. The ${\tt format}$ argument specifies the desired format for standardization. Given the table $T_{6}$ in Figure 2(b), to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%m-%d”.
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $\tt trans\_dict$ to {’Italia’: ’ITA’}.
  • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
Function pool for $\mathtt{Filter}$. We design the Normalizer agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$. This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
Discussions. Our search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools, such as specific data prep functions for various semantic types developed in the dataprep library (Peng et al., 2021).
6. Experiments
6.1. Experimental Setup
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019) and a very recent benchmark TableBench (Wu et al., 2024). The statistics of the datasets are shown in Table 1.
(1) WikiTQ contains complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
(2) TabFact provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Ye et al., 2023) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
(3) TabBench provides an evaluation of table question answering capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation. We evaluate AutoPrep on TabBench-FC and TabBench-NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question. The answers in TabBench could be a string or list of items.
Baselines. We consider the following two categories of baselines:
(1) TQA Methods w/o Data Prep directly generate the answer or utilize programs to extract the answer from the table without considering data prep. We implement four primary TQA baselines.
End2EndQA (End2End) (Chen, 2022) utilizes the in-context learning abilities of LLMs to generate the answer for TQA task based on the supervision of human-designed demonstrations. We implement End2End method with prompt and demonstrations provided by (Cheng et al., 2023).
Chain-of-Thought (CoT) (Wei et al., 2022) prompts LLMs to generate the reasoning process step-by-step before generating the final answer. We implement CoT with the prompt provided by (Chen, 2022).
NL2SQL (Rajkumar et al., 2022) first translates the question into a SQL program and then executes it to get the final answer from the table for the question. We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
NL2Py uses Python code to process and reason over the tables. To construct the prompt for NL2Py, we use TQA instances in NL2SQL prompt and manually write the Python code to process the table and generate the final answer.
Table 1. Statistics of Datasets.
Dataset # Rec. # Row. # Col. Ans. Types
WikiTQ $22,033$ 4$\sim$753 3$\sim$25 string / list (3.05%)
TabFact $2,024$ 5$\sim$47 5$\sim$14 true / false (49.60%)
TableBench $886$ 2$\sim$212 2$\sim$20 string / list (31.49%)
(2) TQA Methods with Data Prep. We also investigate four SOTA TQA methods that consider data prep tasks in their question answering process.
ICL-Prep. uses few-shot in-context learning (ICL) demonstrations to guide LLMs in generating programs for data prep, as shown in Figure 3. Subsequently, we employ NL2SQL to extract answers from the cleaned tables.
Dater (Ye et al., 2023) addresses the TQA task by decomposing the table and question. It first selects relevant columns and rows to obtain a sub-table and then decomposes the origin question into sub-questions. Dater answers these sub-questions based on the sub-tables to generate the final answer. We use code in (cod, 2023b) for implementation.
Binder (Cheng et al., 2023) enhances the NL2SQL method by integrating LLMs into SQL programs. It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone. We utilize the original code provided by (cod, 2023a).
ReAcTable (Zhang et al., 2024) uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs. Once all relevant data is gathered, it asks the LLMs to predict the answer. We run the original code from (cod, 2024b) and keep all settings as default. Notice that the original code does not include prompts for TabFact, we generate it based the WikiTQ prompt.
Chain-of-Table (CoTable) (Wang et al., 2024b) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM. These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM. We implement CoTable using the original code from (cod, 2024a).
Table 2. Improvement of data prep for TQA (the best results are in bold and the second-best are underlined).
Method DeepSeek GPT3.5 Llama3 QWen2.5
WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact
End2End 56.6556.65{56.65}56.65 81.7781.77{81.77}81.77 $52.56$ 71.5471.54{71.54}71.54 58.7258.72{58.72}58.72 81.2781.27{81.27}81.27 60.0160.01{60.01}60.01 81.1781.17{81.17}81.17
    + AutoPrep $63.14$ $\uparrow 6.49$ $82.11$ $\uparrow 0.34$ $61.21$ $\uparrow 8.65$ $71.79$ $\uparrow 0.25$ $61.23$ $\uparrow 2.51$ $84.19$ $\uparrow 2.92$ $63.42$ $\uparrow 3.41$ $83.05$ $\uparrow 1.88$
CoT $54.95$ 82.0282.02{82.02}82.02 53.4853.48{53.48}53.48 $65.37$ $40.75$ 80.9380.93{80.93}80.93 59.6759.67{59.67}59.67 82.3182.31{82.31}82.31
+ AutoPrep $61.12$ $\uparrow 6.17$ $82.26$ $\uparrow 0.24$ $60.01$ $\uparrow 6.53$ $74.36$ $\uparrow 8.99$ $56.01$ $\uparrow 15.26$ $83.65$ $\uparrow 2.72$ $62.02$ $\uparrow 2.35$ $\underline{85.67}$ $\uparrow 3.36$
NL2Py 59.3559.35{59.35}59.35 $68.13$ 53.5953.59{53.59}53.59 66.1566.15{66.15}66.15 $50.12$ $76.24$ $53.02$ $72.63$
+ AutoPrep $\underline{65.86}$ $\uparrow 6.51$ $\underline{87.35}$ $\uparrow 19.22$ $\underline{64.69}$ $\uparrow 11.1$ $\underline{84.83}$ $\uparrow 18.68$ $\underline{62.55}$ $\uparrow 12.43$ $\underline{85.42}$ $\uparrow 9.18$ $\underline{68.65}$ $\uparrow 15.63$ 85.72 $\uparrow 13.09$
NL2SQL $52.83$ $70.21$ $52.90$ $64.71$ 51.8051.80{51.80}51.80 $75.15$ $56.86$ $80.09$
+ AutoPrep 66.09 $\uparrow 13.26$ 87.85 $\uparrow 17.64$ 64.75 $\uparrow 11.85$ 84.19 $\uparrow 19.48$ 63.72 $\uparrow 11.92$ 85.72 $\uparrow 10.57$ 68.72 $\uparrow 11.86$ $85.33$ $\uparrow 5.24$
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the golden answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer, such as pairs like 7.45 and 7.4499, 4 years and 4. Afterwards, we adopt accuracy as our evaluation metric.
Backbone LLMs. We evaluate our method using representative LLMs as backbones. For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613). For open-source LLMs, we choose Llama3 (Dubey et al., 2024) (Llama-3.1-70B-Instruct) and QWen2.5 (Yang et al., 2024) (QWen2.5-72B-Instruct) for evaluation.
Experiment Settings. We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit. For a fair comparison, we set the maximum token input of all methods as 8192. Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
6.2. Improvement of Data Prep for TQA
Exp-1: Impact of question-aware data prep on TQA performance. We integrate AutoPrep into our four TQA baselines w/o data prep, in order to investigate the impact of question-aware data prep on TQA performance. The results are reported in Table 2.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, NL2SQL shows the most substantial gains, achieving an average accuracy improvement of 12.22 on WikiTQ and 13.23 on TabFact across all LLM backbones. Similarly, NL2Py also shows notable improvements in its performance, after being integrated with AutoPrep. This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, End2End and CoT methods also show considerable performance gains. For example, End2End shows an average accuracy gain of 5.27 and 1.35 on WikiTQ and TabFact respectively. These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
Finding 1: The integration of AutoPrep improves the accuracy of TQA methods w/o data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, respectively.
Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
6.3. Data Prep Method Comparison
Table 3. Experimental results of AutoPrep and TQA methods with data prep.
Method DeepSeek GPT3.5 Llama3 QWen2.5
 WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact
ICL-Prep 56.5456.54{56.54}56.54 $80.53$ $55.71$ $73.91$ $50.05$ $75.20$ $57.00$ $80.14$
Dater $48.32$ $83.05$ $52.81$ $72.08$ $43.53$ $74.01$ $58.78$ $79.84$
Binder $56.81$ $82.81$ 56.7456.74{56.74}56.74 79.1779.17{79.17}79.17 $50.51$ $78.16$ $55.43$ 81.7281.72{81.72}81.72
ReAcTable 64.1364.13{64.13}64.13 85.7185.71{85.71}85.71 $51.80$ $72.80$ 58.0158.01{58.01}58.01 80.0080.00{80.00}80.00 60.1560.15{60.15}60.15 $81.67$
CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ $\underline{80.20}$ $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
AutoPrep 66.09 87.85 64.75 84.19 63.72 85.72 68.72 85.33
Exp-2: Comparison of AutoPrep with previous SOTA TQA methods with data prep. We compare AutoPrep with TQA methods that consider data prep tasks in their question answering process, using a single LLM agent. The results are reported in Table 3.
The results clearly demonstrate that AutoPrep achieves the new SOTA performance on both the WikiTQ and TabFact datasets across all LLM backbones. In particular, AutoPrep outperforms ICL-Prep with a significant accuracy improvement by 11.00 and 8.33 on WikiTQ and TabFact on average, respectively. Other baselines, except Dater, generally outperform ICL-Prep by introducing specific technical optimizations to data prep. For instance, ReAcTable applies the ReAct paradigm to iteratively generate data prep operations rather than producing all code at once, and it explicitly considers data prep tasks (e.g., filtering and augmentation). Despite this, AutoPrep still outperforms the second best method by considerable improvement. Specifically, when compared with CoTable, AutoPrep achieves better performance, with an average accuracy improvement by 3.05 and 1.96 on WikiTQ and TabFact, respectively.
These improvements are primarily attributed to our multi-agent framework, which effectively tackles the question-aware data prep challenge. The experimental results indicate that data prep is inherently complex and is hard to be resolved by a single, one-size-fits-all solution; instead, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly. Moreover, AutoPrep covers a broader range of data prep operations than these state-of-the-art TQA methods, thereby addressing gaps (e.g., normalization tasks) not fully explored by previous solutions.
Finding 2: AutoPrep outperforms the the previous SOTA TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact, respectively,
In addition, we find that DeepSeek achieves the better overall performance at lower API costs. Thus, we select DeepSeek as our default backbone LLM for subsequent experiments.
Exp-3: Evaluation of AutoPrep on tables with various sizes. To further investigate AutoPrep, we analyze its performance across tables of varying sizes. While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables. The results are reported in Figure 7.
We observe that all baselines exhibit unstable performance, particularly on larger tables. For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables. Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
(a) Small
(b) Medium
(c) Large
Figure 7. Comparison of AutoPrep and other data prep methods on tables with different sizes on the WikiTQ dataset.
In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes. When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79. The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits. Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
Finding 3: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
6.4. Evaluation on Generalization
Exp-4: Evaluating generalization on unseen datasets. As discussed previously, we evaluate the generalization capabilities of AutoPrep on two TabBench datasets. Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods. Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capability of our method. The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
Finding 4: Utilizing tool-augmented execution enhances the generalization capabilities of AutoPrep on unseen datasets.
(a) Numerical Reasoning
(b) Fact Checking
Figure 8. Evaluating Generalization on the TabBench dataset.
6.5. Ablation Studies
Exp-5: Evaluation on the Planner agent. We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 4(a).
We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively. This indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using Direct Prompting.
Finding 5: Our proposed Chain-of-Clauses method is more effective in suggesting high-level operations.
(a) Case one
(b) Case two
Figure 9. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
Exp-6: Evaluation on the Programmer Agents. We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
As illustrated in Table 4(b), our proposed tool-augmented method achieves better performance compared with the code generation method. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. We also record the error ratio of these two methods, as shown in Table 4(b). The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
Finding 6: Our tool-augmented method performs better in both accuracy and cost, with fewer program errors.
Table 4. Experimental Results of Ablation Studies.
(a) Evaluation on the Planner agent.
Method  WikiTQ  TabFact  TabBench
Direct Prompting $58.17$ $83.35$ $44.83$
Chain-of-Clauses 66.09 87.85 50.10
(b) Evaluation on the Programmer Agents.
Method Metric  WikiTQ  TabFact  TabBench
Code Generation Acc $\uparrow$ $62.82$ $81.97$ $47.26$
Err $\downarrow$ $4.51\%$ $4.50\%$ $4.73\%$
Tool Augmented Acc $\uparrow$ 66.09 87.85 50.10
Err $\downarrow$ 1.50% 0.05% 1.01%
(c) Contribution of Each Programmer Agent in AutoPrep.
Method  WikiTQ  TabFact  TabBench
AutoPrep 66.09 87.85 50.10
- Filter $62.78$ ($-3.31$) $84.98$ ($-2.87$) $47.67$ ($-2.43$)
- Augmenter $61.79$ ${(-4.30)}$ $85.67$ ($-2.18$) $45.44$ ($-4.66$)
- Normalizer $62.02$ ($-4.07$) $83.79$ ${(-4.06)}$ $41.58$ ${(-8.52)}$
Exp-7: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
As shown in Table 4(c), for WikiTQ, without column augmentation, the accuracy drops the most (4.30). For TabFact and TabBench, the Normalizer matters the most with an accuracy drop by 4.06 and 8.52. This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while normalization is a primary issue in TabFact and TabBench. Moreover, for all datasets, each agent plays an essential role in data preparation for TQA, which brings accuracy improvement by at least 2.43, 2.18 and 4.06.
Finding 7: Solving three types of data preparation tasks in TQA methods in a specialized manner improves the downstream TQA performance.
6.6. Case Studies
This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
To answer the question in Figure 9(a), CoTable generates an operator chain with an “add_col” operator, which is used to generate a new column Tol based on the division results of two existing columns Expense and Percent. This operator directly prompts an LLM to output a list containing all values of the new column, which is a challenging task for previous LLMs. Thus, CoTable generates a new column with two wrong values which leads to a wrong answer. For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation. Thus, ReAcTable also fails to augment the Tol column. AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Augment}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol. These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function. For example, we call $\mathtt{to\mbox{-}numerical}$ function two times to cast Expense and Percent to numerical values. All physical operators are executed to process the original table and output a prepared table. Finally, an Analyzer agent based on NL2SQL method generates a SQL query to extract the correct answer “4214.17”.
Figure 9(b) illustrates a TQA instance involving large tables and inconsistent values. AutoPrep solves this by splitting the TQA question into two phases, i.e., data prep and data analysis. It first generates logical operators including two $\mathtt{Normalize}$ and one $\mathtt{Filter}$. Next, the Programmer agents implement them with corresponding physical operators, which are executed to produce a prepared table. Based on this, the Analyzer agent extracts the final answer “8”.
7. Related Work
Tabular Question Answering. Most of the SOTA solutions for TQA rely on LLMs (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables. Specifically, direct Prompting approaches (Chen, 2022; Wei et al., 2022), which prompt an LLM to directly generate answers through next token prediction, may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024). Code Generation approaches (Rajkumar et al., 2022) leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and adaptability.
Previous TQA methods, like Dater (Ye et al., 2023), Binder (Cheng et al., 2023), CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), also consider data prep in their question answering process. Specifically, Dater (Ye et al., 2023) prompts LLMs to select relevant columns related to answering the question, targeted at solving filtering tasks. Binder (Cheng et al., 2023) proposes to integrate SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks. Similarly, CoTable (Wang et al., 2024b) addresses the filtering tasks and augmentation tasks by designing operators which are implemented by LLM completion. ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the augmentation and filtering tasks.
However, previous TQA methods have the following limitations when handling data prep. First, all these methods do not address the cell-level normalization tasks, which account for the most significant error types, as indicated in Figure 1. Second, all these methods utilize a single LLM agent for both data prep and answer reasoning. Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
Traditional Data Preparation. Data prep techniques have been widely adopted in various tasks. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations. Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
In this paper, we propose AutoPrep to generate question-specific data prep operations for TQA tasks. The key difference between these methods and AutoPrep is that the data prep operations generated by AutoPrep are tailored to specific user questions and customized for the tables, as different questions may have different data prep requirements, even for the same table.
LLMs-based Multi-Agent Framework. A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a). Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024). AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Tabular Question Answering.
While AutoTQA improves table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
8. Conclusion
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and report error messages. We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Additionally, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ, TabFact and TabBench demonstrate that: (1) AutoPrep achieves SOTA performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the performance of TQA methods.